# Field of View Enhanced Signal Dependent Binauralization with Mixture of Experts Framework for Continuous Source Motion

### v1

@@ Abstract
We propose a novel mixture of experts framework for field-of-view enhancement in binaural signal matching. Our approach enables dynamic spatial audio rendering that adapts to continuous talker motion, allowing users to emphasize or suppress sounds from selected directions while preserving natural binaural cues. Unlike traditional methods that rely on explicit direction-of-arrival estimation or operate in the Ambisonics domain, our signal-dependent framework combines multiple binaural filters in an online manner using implicit localization. This allows for real-time tracking and enhancement of moving sound sources, supporting applications such as speech focus, noise reduction, and world-locked audio in augmented and virtual reality. The method is agnostic to array geometry offering a flexible solution for spatial audio capture and personalized playback in next-generation consumer audio devices.
@@ 1 Introduction
Consumer audio capture devices are increasingly designed as wearable technologies. Among these, headworn microphone arrays have gained significant attention for capturing sound fields and enabling binaural rendering. A key use case arises when the user wishes to re-experience the recording in a way that matches how it originally sounded. This places importance on downstream processing methods that preserve the auditory cues present at the time of capture [1]. The process of filtering and summing the microphone array signals to reproduce the binaural cues at the left and right ears of the user has been referred to as end-to-end magnitude least squares (eMagLS) [2] or binaural signal matching (BSM) [3]. An alternative method to render binaural signals relies on direction-of-arrival (DOA) estimation and beamforming to extract direct signal components and renders direct and reverberant sound field components separately [4]. It is desirable to give the user additional control over properties of the rendered audio. Such control may take the form of speech enhancement, noise reduction, support for world-locked audio where playback adapts to user movement, or directional enhancement where sounds from selected directions or from the user’s gaze are emphasized [5, 6, 7]. This paper develops a signal-dependent framework for the latter task, which we refer to as field-of-view enhancement (FoVE) [8]. Moreover, we propose a mixture of experts algorithm that is able to combine the estimates of numerous signal-dependent binaural signal matching filters in an online manner with implicit localization instead of relying on traditional direction of arrival estimators. This allows the model to track a continuous talker without assumptions of stationarity. Figure 1 depicts the system diagram for the proposed method.
@@ 2 Signal Model
In the next section, we review both signal-dependent and independent variants of binaural signal matching. At high frequencies, the magnitude-LS variant replaces the complex LS objective while using the same $\hat{\mathbf{R}}_{s}$ construction.
@@ 3.1 Signal-Independent Binaural Signal Matching
Signal-independent BSM aims to design a linear filter that maps the microphone array signals to binaural signals at the user’s ears. The design does not depend on a specific source signal but instead assumes a diffuse sound field. This corresponds to energy being uniformly distributed across all directions of arrival. The least-squares optimal binaural filter weights $\mathbf{c}^{\text{BSM}}\in\mathbb{C}^{2\times N_{m}}$ solve,
@@ 3.2.1 COMPASS-BSM (COM)
COMPASS is a parametric method used to render spatial audio scenes [4]. The directions of arrival of sources are estimated first. Given those directions, the received soundfield is decomposed into direct and residual parts, where the direct part is obtained as the linearly constrained minimum variance [9] (LCMV) estimate
@@ 3.2.2 Directional BSM (d-BSM)
Form a source covariance that embeds the estimated direct component and a diffuse reverberant component[10]:
@@
where $\hat{\sigma}_{r}^{2}$ is obtained from the residual consistent with the diffuse assumption in the paper. Let $\mathbf{A}=[\,\mathbf{A}_{d}\ \mathbf{A}_{r}\,]$ stack the direct and reverberant steering matrices and let $\mathbf{h}$ denote the ear-specific HRTF vector. The signal-dependent BSM filter is
@@ 3.3 Proposed Method
$\mathbf{c}^{MoE}[t,f]$ $=\sum_{q=1}^{Q}\alpha_{q}[t,f]\,\mathbf{c}_{q}[t,f],$ (18)
$\hat{\mathbf{p}}[t,f]$ $=\big{(}\mathbf{c}^{\text{MoE}}\big{)}^{H}[t,f]\,\mathbf{X}[t,f].$ (19)
Intuitively, this is like an implicit DOA estimator that relies on the assumption that the beamformer with the lowest residual likely contains the direct source.
@@ 3.3.1 Regret Bound
This guarantee implies that, asymptotically, the mixture performs nearly as well as the best expert chosen with full hindsight, while retaining the flexibility to adapt to time-varying acoustic scenes. Note that the average regret $R_{T}/T$ goes to zero as time approaches infinity. [13]
@@ 4.1 Gain Control
where $\gamma\in[0,1]$ is a user-selected gain parameter. Note, that for COMPASS we can add gain $\geq 1$ in the FoV directions that are rendered directly through the HRTF as shown in equation 6.
@@ 4.2 Distortion Control
$\min_{\mathbf{c}}\;\big{(}\mathbf{c}\mathbf{A}-\mathbf{H}\big{)}\mathbf{D}\big{(}\mathbf{c}\mathbf{A}-\mathbf{H}\big{)}^{T},$ (24)
@@
In directional BSM (d-BSM), the distortion weights enter the covariance matrices. The weighted source covariance is
@@
$\mathbf{c}=\mathbf{H}^{T}\,\tilde{\mathbf{R}}_{s}\,\mathbf{A}^{H}\big{(}\mathbf{A}\,\tilde{\mathbf{R}}_{s}\,\mathbf{A}^{H}+\mathbf{R}_{n}\big{)}^{-1},$ (26)
@@ 4.3 Proposed FoVE via Mixture of Experts
The mixture of experts framework described earlier can be directly extended to field of view enhancement. Each expert corresponds to a binaural filter computed under a FoV-aware formulation, and the adaptive blending is performed in the same manner as before. In practice, the proposed method can be made FoV-enhanced simply by constructing the expert filters using any of the FoV-aware versions of BSM, COMPASS-BSM, or d-BSM. The resulting mixture the emphasis on user-selected FoV regions. Figure 2 depicts the direction dependent gain from varying the two parameters $\gamma$ and $\delta$ for signal independent beamforming. The same results extends to each of the signal-dependent methods relative the weights without FoVE. Note, that we elect to use magnitude least-squares optimization or MagLS. [14, 15, 2]
Fig. 2: The direction dependent gain from field of view in the frontal direction. The plot on left shows the effect of varying $\gamma$ with a fixed value of $\delta$ while the second plot shows the effect of varying $\delta$ with fixed value of $\gamma$
@@ 5.1 Simulation
A continuous motion simulation is created in pyroomacoustics [16] in a room of size [8m, 8m, 5m] with RT60 $\approx$ 200ms. A 4-microphone array is simulated with the reference center at [4m, 4m, 2m]. All signals are sampled synchronously at 48 kHz and the talker speech is derived from the EARS dataset [17]. One moving talker is simulated with a 6-degree spacing in azimuthal direction. The talker is initialized to position [7m, 4m, 2m] in the frontal direction of the microphone array. The talker covers every 6 degrees in 167ms, which equates to a speed $\approx$ 2 $m/s$. Figure 3 demonstrates the efficacy of the proposed residual based loss, that is clearly able to track the talker in the volume. Additionally, we have depicted the value of blend weights for the set of BSM filters, that shows that in continuous motion the algorithm prefers to blend different filters resulting in perceptually relevant directions being rendered correctly resulting in accurate interaural cues at each time.
Fig. 3: The argmin of the residual energy at each time and the corresponding blend weights for the grid of beamformers for the simulation
@@ 5.2 Experimental
The simulation experiment is recreated in a real-world environment with a 4 microphone headworn microphone array sampled at 48 kHz. One talker was asked to start in front of the listener and walked in a counter-clockwise direction similar to the simulation experiment. Figure 4 demonstrates the effectiveness of the proposed method in tracking the talker as they walk in the environment. Depicted in the same figure is the blend weights of the different candidate binauralization filters over time. This clearly demonstrates that we are able to render the source with the correct binaural cues as they move through the environment especially with MoE module that would render directly using the HRTFs.
Fig. 4: The argmin of the residual energy at each time and the corresponding blend weights for the grid of beamformers in the experiment performed
@@ 6 Conclusion
In this work, a novel mixture of experts framework is theorized for binauralization. The proposed framework extends previous work in signal-dependent binauralization to scenarios with continuous motion and for adjustable field of view enhancement. Our results demonstrate that the framework is not only effective but highly modular, giving a user fidelity over the reproduced acoustic scene. Our approach leverages a softmax function to seamlessly blend outputs from numerous BSM filters, each tailored to distinct direction. This modularity enables real-time adaptation to listener movement and environmental changes, ensuring consistent spatial audio quality. Future work will focus on neural beamspace projection.

### v3

# Mixture-of-Experts Framework for Field-of-View Enhanced Signal-Dependent Binauralization of moving talkers

@@ Abstract
We propose a novel mixture of experts framework for field-of-view enhancement in binaural signal matching. Our approach enables dynamic spatial audio rendering that adapts to source motion, allowing users to emphasize or suppress sounds from selected directions while preserving natural binaural cues. Unlike traditional methods that rely on explicit direction-of-arrival estimation, our signal-dependent framework combines multiple binaural filters in an online manner using implicit localization. This allows for real-time tracking and enhancement of moving sound sources, supporting applications such as speech focus, noise reduction, and world-locked audio in augmented and virtual reality. The method is agnostic to array geometry and offers a flexible solution for spatial audio capture and playback in next-generation consumer audio devices.
@@ 1 Introduction
Consumer audio capture devices are increasingly designed as wearable technologies. Among these, headworn microphone arrays have gained significant attention for capturing sound fields and enabling binaural rendering. A key use case arises when the user wishes to re-experience the recording in a way that matches how it originally sounded. This places importance on downstream processing methods that preserve the auditory cues present at the time of capture [1]. The process of filtering and summing the microphone array signals to reproduce the binaural cues at the left and right ears of the user has been referred to as end-to-end magnitude least squares (eMagLS) [2] or binaural signal matching (BSM) [3]. An alternative method to render binaural signals relies on direction-of-arrival (DOA) estimation and beamforming to extract direct signal components and renders direct and reverberant sound field components separately [4].
It is desirable to give the user additional control over properties of the rendered audio. Such control may take the form of speech enhancement, noise reduction, support for world-locked audio where playback adapts to user movement, or directional enhancement where sounds from selected directions or from the user’s gaze are emphasized [5, 6, 7]. This paper develops a signal-dependent framework for the latter task, which we refer to as field-of-view enhancement (FoVE) [8]. Moreover, we propose a mixture of experts algorithm that is able to combine the estimates of numerous signal-dependent binaural signal matching filters in an online manner with implicit localization instead of relying on traditional direction of arrival estimators. This allows the model to track a continuous talker without assumptions of stationarity. Fig. 1 depicts the system diagram for the proposed method.
@@ 2 Signal Model
In the next section, we review both signal-dependent and independent variants of binaural signal matching. At high frequencies, the magnitude-LS variant replaces the complex LS objective while using the same $\hat{\mathbf{R}}_{s}$ construction [9, 2, 3].
@@ 3.1 Signal-Independent Binaural Signal Matching
Signal-independent BSM aims to design a linear filter that maps the microphone array signals to binaural signals at the user’s ears [2, 3]. The design does not depend on a specific source signal but instead assumes a diffuse sound field. This corresponds to energy being uniformly distributed across all directions of arrival. The least-squares optimal binaural filter weights $\mathbf{c}^{\text{BSM}}\in\mathbb{C}^{2\times N_{m}}$ solve,
@@ 3.2.1 COMPASS-BSM (COM)
COMPASS is a parametric method used to render spatial audio scenes [4]. The directions of arrival of sources are estimated first. Given those directions, the received soundfield is decomposed into direct and residual parts, where the direct part is obtained as the linearly constrained minimum variance (LCMV) [10] estimate
@@ 3.2.2 Directional BSM (d-BSM)
Signal-dependent BSM (d-BSM) utilizes a source covariance that embeds the estimated direct component, which is estimated as in COMPASS, and a diffuse reverberant component[11]:
@@
where $\hat{\sigma}_{r}^{2}$ is obtained from the residual consistent with the diffuse assumption in the paper. Let $\mathbf{A}=[\,\mathbf{A}_{d}\ \mathbf{A}_{r}\,]$ stack the direct and reverberant steering matrices and The signal-dependent BSM filter is
@@ 3.3 Proposed Method
$\mathbf{c}^{\text{MoE}}[t,f]$ $=\sum_{q=1}^{Q}\alpha_{q}[t,f]\,\mathbf{c}_{q}[t,f],$ (18)
$\hat{\mathbf{p}}[t,f]$ $=\big(\mathbf{c}^{\text{MoE}}[t,f]\big)^{H}\,\mathbf{X}[t,f].$ (19)
Intuitively, the method behaves like an implicit DOA estimator that relies on the assumption that the beamformer with the lowest residual likely contains the direct source.
@@ 3.3.1 Regret Bound
This guarantee implies that, asymptotically, the mixture performs nearly as well as the best expert chosen with full hindsight, while retaining the flexibility to adapt to time-varying acoustic scenes. Note that the average regret $R_{T}/T$ goes to zero as time approaches infinity [14].
@@ 4.1 Gain Control
where $\gamma\in[0,1]$ is a user-selected gain parameter. Note, that for COMPASS we can add gain $\geq 1$ in the FoV directions that are rendered directly through the HRTF as shown in (6).
@@ 4.2 Distortion Control
$\min_{\mathbf{c}}\;\big(\mathbf{c}\mathbf{A}-\mathbf{H}\big)\mathbf{D}\big(\mathbf{c}\mathbf{A}-\mathbf{H}\big)^{H},$ (24)
@@
In d-BSM, the distortion weights enter the covariance matrices. The weighted source covariance is
@@
$\mathbf{c}=\mathbf{H}^{T}\,\tilde{\mathbf{R}}_{s}\,\mathbf{A}^{H}\big(\mathbf{A}\,\tilde{\mathbf{R}}_{s}\,\mathbf{A}^{H}+\mathbf{R}_{n}\big)^{-1}.$ (26)
@@ 4.3 Proposed FoVE via Mixture of Experts
The mixture of experts framework described earlier can be directly extended to field-of-view enhancement simply by constructing the expert filters using any of the FoV-aware versions of BSM, COMPASS-BSM, or d-BSM. As shown in Fig. 2, the resulting mixture successfully places emphasis on user-selected FoV regions.
Fig. 2: The direction dependent gain from field of view in the frontal direction. The plot on left shows the effect of varying $\gamma$ with a fixed value of $\delta$ while the second plot shows the effect of varying $\delta$ with fixed value of $\gamma$.
@@ 5.1 Simulation
A continuous motion simulation is performed in pyroomacoustics [15] within an [8 m, 8 m, 5 m] room (RT60 $\approx$ 200 ms). A 4-microphone array centered at [4 m, 4 m, 2 m] records speech from the EARS dataset [16], sampled at 48 kHz. One talker, initialized at [7 m, 4 m, 2 m] in front of the array, moves in $6^{\circ}$ azimuth steps, covering each step in 167 ms ($\approx$ 2 m/s). Fig. 3 (top) shows that the proposed residual-based loss accurately tracks the talker’s motion. The blend weights of the BSM filters further indicate that, under continuous motion, the algorithm adaptively combines filters to render perceptually relevant directions.
Fig. 3: The argmin of the residual energy at each time and the corresponding blend weights for the grid of beamformers for the simulation (top) and the measurement (bottom).
@@ 5.2 Experimental
The simulation experiment is recreated in a real-world environment with a 4-microphone head-worn microphone array. One talker was asked to start in front of the listener and walk in a counter-clockwise direction, similar to the simulation experiment. Fig. 3 (bottom) demonstrates the effectiveness of the proposed method in tracking the talker as they walk in the environment. This indicates that, also under real-world conditions, the proposed method is able to render the direct source component with the correct binaural cues as they move through the environment, especially with the MoE module that directly utilizes the HRTFs. Shown in figure  4 is the error in the ITD and ILD from each of the candidate algorithms. This corresponds to the error is rendering the residual sound field and demonstrates that we are able to match the interaural cues from BSM while providing the user with additional fidelity over the recorded scene. Notably, we could enhance the moving source dynamically over time using the proposed framework with a set of FoVE filters designed to enhance each direction.
Fig. 4: The error in ITD and ILD over time for the grid of directions considered on average. Results demonstrate that we can manipulate the FoV of the acoustic scene with equally accurate binaural cues as previous methods.
@@ 6 Conclusion
In this work, a novel mixture of experts framework is theorized for binauralization. The proposed framework extends previous work in signal-dependent binauralization to scenarios with continuous motion and for adjustable field-of-view enhancement. Our results demonstrate that the framework is not only effective but highly modular, so that it can be extended to incorporate new advances in the field. For example, future work will focus on utilizing a neural beamspace projection.